OpenStudy (anonymous):

which of the following sets do not form a right triangle? 6,8,10 8,15,17 7,24,26 5,12,13

OpenStudy (anonymous):

So you know about pythagoras theorem?

OpenStudy (anonymous):

no not really?

OpenStudy (anonymous):

Ok the thing is; in any right triangle, the square of the bigger side equals sum of the squares of smaller sides. so with 6,8,10, 6^2 + 8^2 = 36 + 64 = 100 which is 10^2. so u can try with the other 3 and see which one doesn't fulfill this condition.

OpenStudy (anonymous):

The third one does not have a right angle because 7^2 + 24^2 = 625. Then the square root of 625 is 25.